Solve each system by the substitution method (please):
4x - y = 100
0.05x - 0.06y = -32
Thank you ^_^

Answer by mananth(16946)

4x - y = 100
y=4x-100
0.05x - 0.06y = -32
multiply by 100
5x-6y=-3200
substitute y in the equation
5x-6(4x-100)=-3200
5x-24x+600=-3200
-19x=-3800
/-19
x=200
plug value of y in 4x-y=100
800-y=100
y=800-100
y=700
m.ananth@hotmail.ca